OpenStudy (imnotgoodatmath):

Which expression is equivalent to 11^-2?

OpenStudy (anonymous):

1/121

OpenStudy (haseeb96):

1/11^2 = 11 ^-2

OpenStudy (anonymous):

give your options..???

OpenStudy (anonymous):

@ImNotGoodAtMath

OpenStudy (imnotgoodatmath):

oh, sorry, playing dark souls 2 lol

OpenStudy (haseeb96):

yes what are yur opyions? but if there is u are asking about which expression equal to that so answer is 1/11 ^2 because when we reverse it then we will get same equation as written in the above given question

OpenStudy (imnotgoodatmath):

eleven times negative two one divided by the quantity eleven times eleven one divided by the quantity eleven times two times two negative eleven times two times two

OpenStudy (anonymous):

its good home work ..:P :P :P

OpenStudy (haseeb96):

C is the correct i wrote that in mathematical form 1/11^2 = 1/11 into 11

OpenStudy (anonymous):

sorey..its option B.... one diveded by 11 times 11...

OpenStudy (haseeb96):

yes B is the correct answer C is not the correct answer because i have mistaken because i was hurry in solving the question